# An Iterative Bayesian Approach for System Identification based on Linear Gaussian Models

### v2

@@ Abstract
We tackle the problem of system identification, where we select inputs, observe the corresponding outputs from the true system, and optimize the parameters of our model to best fit the data. We propose a flexible and computationally tractable methodology that is compatible with any system and parametric family of models. Our approach only requires input-output data from the system and first-order information from the model with respect to the parameters. Our algorithm consists of two modules. First, we formulate the problem of system identification from a Bayesian perspective and use a linear Gaussian model approximation to iteratively optimize the model’s parameters. In each iteration, we propose to use the input-output data to tune the covariance of the linear Gaussian model. This statistically calibrates the approach. Secondly, we define a Gaussian-based uncertainty measure for the model parameters, which we can then minimize with respect to the next selected input. We test our method with linear and nonlinear dynamics.
@@ I Introduction
In this work, we are interested in the problem of fitting a model to a system and choosing informative inputs to improve the model. Our formulation tackles the general problem of model-fitting and dataset collection, which is known as system identification. Obtaining a good model of the system is a key requirement in various engineering tasks. System identification is usually the first step in a larger pipeline that involves the control of the underlying system, e.g., with model predictive control [1]. Fitting model parameters to data can be challenging when the model is a complicated function of the parameters. In addition, Ott et al. have shown that selecting control inputs based on the current model can improve the system identification process [2]. Informative input design can reduce the data gathering requirements for the purpose of modeling the system, which can be costly in real-world applications. This can be particularly useful in high-dimensional problems, where finding informative directions to perturb the system can be challenging.
There are four basic elements in the problem of system identification. We need to select a parametric family of models for the underlying system. The family must be expressive enough to capture the true system, but also simple enough to optimize over. Our parameter estimate is improved once more data becomes available and we must also choose the optimization approach that updates the model parameters. A vast literature exists on methods for estimating the model parameters [3]. Furthermore, we need an uncertainty measure to minimize over the next input. This allows us to select inputs that best inform our choice of model parameters. Various objectives based on the Fisher information matrix can be minimized over the input [4]. This optimization problem must also be accompanied by an optimization approach.
@@
  • Can we design a method for system identification that is computationally tractable, compatible with any parametric model family and able to suggest whether a parametric model family is suitable for a given system?
Our approach leverages first-order information to construct a linear Gaussian model for the output given the input, as a function of the model parameters. Based on the first-order Taylor expansion, we solve successive convex problems, corresponding to maximum likelihood estimation (MLE), to update the model parameters. This simplifies the model update step discussed in the previous paragraph, which becomes computationally tractable. We also propose a procedure to tune the covariance associated with the linear Gaussian model. This allows us to statistically calibrate our model and judge the suitability of the parametric model family. Finally, using the linear Gaussian model from the first-order Taylor expansion, we define an uncertainty measure that is accurate in the linear Gaussian case for the input design.
@@ II Prior Work
The problem of system identification has been well studied. Early research on the optimal selection of inputs dates back to the 1960s [5, 6] and highlights the importance of the problem for the identification of systems. Recent advances in computational capabilities and learning-based methods have once again brought attention to the problem [7, 8]. This section summarizes the prior work on parameter estimation for system identification and informative input design.
Estimating the model parameters using input-output data is usually formulated as a regression problem [9]. A regression problem can be equivalently viewed as MLE under an appropriate distribution. For example, least squares regression corresponds to MLE with a Gaussian distribution. When estimating the model parameters, it is common to assume the true system belongs to the parametric model family [3, 4, 10]. Then, under regularity conditions on the distribution of the input or the structure of the system, the parameter estimate is concentrated around the true system parameter [3]. Our work deals with the problem of system identification when the true system does not necessarily belong to the selected parametric model family. In this case, it is critical to obtain a good fit but also a sense of the suitability of the selected model family.
@@
In various applications, the system contains modes that are not properly described by a linear model. In these cases, we must deal with nonlinear models and nonlinear system identification. Early results focused on specific system dynamics for which the input design problem can either be relaxed using linear matrix inequalities [13] or techniques from linear systems can be applied to design informative step inputs [14]. In certain instances, the nonlinearity can be separated from the linear part of the model in what is known as structured nonlinear system identification [4]. There is consensus among researchers that nonlinear system identification is much more difficult than linear system identification [15]. A simple justification can be made by examining the difficulty of solving a nonlinear regression problem. The problem of informative input design for nonlinear dynamical systems remains a large unexplored area of research [4]. Our approach deals with the identification of nonlinear dynamics.
We examine methods for nonlinear system identification that are relevant to our approach. Goodwin presented an algorithm for informative input design in the general nonlinear case that is based on linearization of the parametric model [16]. Our work adds to this work in an important way: our method updates the estimate of the model parameters in a way that respects the linearization and is based on convex optimization, which makes it computationally fast. A similar work was done by Babar and Baglietto [17], who propose an extended Kalman filter for the joint state and parameter estimation of the dynamical system. Model predictive control techniques are used to select the next input. We improve upon their work by adding a trust-region constraint to the update of the model parameters after the linearization and by proposing a modified optimization objective for the selection of the next input that does not rely on sampling. In contrast with both works [16, 17], we tune the noise covariance at every iteration. This statistically calibrates our approach and provides information on the suitability of the selected model family for the system in question. We show how the approach can fail without tuning the covariance.
@@ III Problem Formulation
where $x\in\mathbb{R}^{d_{x}},y\in\mathbb{R}^{d_{y}}$ and $g:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{y}}$. If $g$ represents dynamics, the input is usually the current state and control, which we abstract as $x$, while the next state is abstracted as $y$. We refer to $x$ as the input and $y$ as the output. We treat system $g$ as a black-box. We can only supply an input $x$ and provide an output $y$.
We choose a family of functions $f(\cdot;\theta)$, parameterized by $\theta\in\mathbb{R}^{q}$. At every iteration, our goal is to compute:
@@
    $\hat{\theta}$ such that $g(\cdot)\approx f(\cdot;\hat{\theta})$ using the available data,
@@ IV Proposed Method
We first outline our Bayesian approach for improving the current estimate of the model parameters. We then use this Bayesian formulation in order to define an optimization problem for the next input. The components presented so far can be found in the literature [16], with the exception of the trust region in the estimate improvement step and the ability to choose the noise probability density function. Finally, we discuss the fine-tuning, which is critical in the convergence of the method and constitutes our major contribution.
@@ IV-A Parameter Estimation
The collected data satisfy eq. 2. Because the current model $f(\cdot;\hat{\theta})$ might not match the true system $g(\cdot)$ exactly, it holds that
@@
for some noise realization $\epsilon_{i}$. We choose the underlying noise model as $\epsilon_{i}\sim\mathcal{N}(0,\Sigma)$ for some covariance $\Sigma$. Tuning this covariance is critical in the performance of the algorithm. Assuming conditionally independent measurements, the posterior for the hidden parameter is given by
@@
where we choose the prior $\theta\sim\mathcal{N}(\theta_{\mathrm{prior}},\Sigma_{\mathrm{prior}})$.
The maximum a posteriori (MAP) estimate for the parameter under these assumptions is given by
@@
$=\operatorname*{arg\;min}_{\theta}\ \dfrac{1}{2}\lVert\theta-\theta_{\mathrm{prior}}\rVert^{2}_{\Sigma_{\mathrm{prior}}^{-1}}+$
$\qquad\dfrac{1}{2}\sum_{i=1}^{n}\lVert y_{i}-f(x_{i};\theta)\rVert_{\Sigma^{-1}}^{2}.$
In general, this can be a very difficult problem to solve because it depends on the structure of the parametric family $f$ with respect to the parameter $\theta$. We opt for an approximate treatment of the optimization problem that lends itself to a convex solver [18].
@@
$f(x;\theta)$ $\approx f(x;\hat{\theta})+\dfrac{\partial f}{\partial\theta}|_{(x;\hat{\theta})}\left(\theta-\hat{\theta}\right)$ (6)
@@
$b(x;\hat{\theta})$ $=f(x;\hat{\theta})-\dfrac{\partial f}{\partial\theta}|_{(x;\hat{\theta})}\hat{\theta},$
$C(x;\hat{\theta})$ $=\dfrac{\partial f}{\partial\theta}|_{(x;\hat{\theta})}.$
@@
$\hat{\theta}_{+}\approx\operatorname*{arg\;min}_{\lVert\theta-\hat{\theta}\rVert\leq\delta}\ \dfrac{1}{2}\lVert\theta-\theta_{\mathrm{prior}}\rVert^{2}_{\Sigma_{\mathrm{prior}}^{-1}}$ (8)
$+\dfrac{1}{2}\sum_{i=1}^{n}\lVert\underbrace{y_{i}-b(x_{i};\hat{\theta})}_{\tilde{y}_{i}}-C(x_{i};\hat{\theta})\theta\rVert_{\Sigma^{-1}}^{2}.$
This is a convex problem with respect to $\theta$ and can easily be handled by a computational framework for convex optimization [19]. The constraint $\lVert\theta-\hat{\theta}\rVert\leq\delta$ guarantees that we search for the next model parameter in the region where the Taylor series for $f(\cdot;\theta)$ is sufficiently close to $f(\cdot;\theta)$. It also does not allow our next estimate to differ much relative to the current one. This is justified because the next input-output pair should not change our estimate drastically. Placing too much trust on any one datapoint can lead to excessively large steps or premature convergence [20].
We can replace Gaussian densities with any other log concave density and still obtain a convex problem. The used density for $\epsilon_{i}$ can be changed in each estimation step, based on the realized errors $\{y_{i}-f(x_{i};\hat{\theta})\}_{i=1}^{n}$. This allows us to select among a family of parameter estimates. More details on this will be given in section IV-C.
@@ IV-B Informative Input Design
After obtaining $\hat{\theta}_{+}$, our goal is to select $x_{n+1}$. The goal of selecting $x_{n+1}$ is to reduce our uncertainty in the posterior estimate of $\theta$. Assuming that we are given a new data-point $(x_{n+1},y_{n+1})$ and under completely linear Gaussian assumptions, the posterior inverse covariance is given by the Hessian of the objective of (8) but evaluated using the most up-to-date MAP estimate $\hat{\theta}_{+}$, i.e,
$\Sigma_{\mathrm{posterior}}^{-1}=\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma^{-1}C(x_{i};\hat{\theta}_{+}).$ (9)
It has been shown that such an expression for the covariance can be accurate in the nonlinear case, although it is exact only in the linear regime [21]. Eq. (9) is an approximate Fisher information matrix, where a prior is added and the expectation over the data is not used. We note that it only depends on the inputs $x_{1},\dots,x_{n+1}$ and not on the corresponding outputs.
@@
$\mathcal{M}\left(\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma^{-1}C(x_{i};\hat{\theta}_{+})\right),$ (10)
@@
$\operatorname*{arg\;max}_{x_{n+1}}$ $\mathcal{M}\left(\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma^{-1}C(x_{i};\hat{\theta}_{+})\right)$ (11)
@@
In order to solve eq. 11 with gradient descent methods, we can replace the indicators with smooth functions. By increasing parameter $\lambda\geq 0$, we can move the search towards feasible inputs $x_{n+1}$. This is an example of a penalty method [20]. As a receding horizon extension, we can easily extend eq. 11 to optimize over the next $k$ inputs, $x_{n+1},\dots,x_{n+k}$.
IV-C Finetuning
We present the finetuning approach for the covariance $\Sigma$ that is critical for the convergence of our algorithm. We also propose a way to tune $\delta$.
At the current iteration , we have access to the dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$. After obtaining the new parameter estimate $\hat{\theta}_{+}$, we can compute the estimated noise realizations $\hat{\epsilon}_{i}$ by comparing the true output to the expected output, i.e., for $i=1,\dots,n$
$\hat{\epsilon}_{i}=y_{i}-f(x_{i};\hat{\theta}_{+}).$ (12)
Using these samples, we can compute a covariance for the errors caused by the mismatch between $g(\cdot)$ and $f(\cdot;\hat{\theta}_{+})$:
$\Sigma_{\mathrm{model\ error}}=\dfrac{1}{n}\sum_{i=1}^{n}\hat{\epsilon}_{i}\hat{\epsilon}_{i}^{\top}.$ (13)
@@ IV-C Finetuning
$y_{i}-b(x_{i};\hat{\theta})-C(x_{i};\hat{\theta})\theta$ (14)
@@
$\hat{\epsilon}_{i}+\underbrace{f(x_{i};\hat{\theta}_{+})-\left(b(x_{i};\hat{\theta})+C(x_{i};\hat{\theta})\hat{\theta}_{+}\right)}_{\bar{\epsilon}_{i}}.$
This implies that the covariance $\Sigma$ should be set using the sample covariance of $\{\hat{\epsilon}_{i}+\bar{\epsilon}_{i}\}_{i=1}^{n}$:
$\Sigma\leftarrow\dfrac{1}{n}\sum_{i=1}^{n}\left(\hat{\epsilon}_{i}+\bar{\epsilon}_{i}\right)\left(\hat{\epsilon}_{i}+\bar{\epsilon}_{i}\right)^{\top},$ (16)
where $\bar{\epsilon}_{i}$ is the error between the value from the model and the value from the Taylor expansion for the $i$-th output.
For setting the trust region $\delta$, we can use a simple heuristic. If the average linearization error size $n^{-1}\sum_{i=1}^{n}\lVert\bar{\epsilon}_{i}\rVert$ becomes comparable to the average absolute model error size $n^{-1}\sum_{i=1}^{n}\lVert\hat{\epsilon}_{i}\rVert$, we reduce $\delta$. This procedure guarantees that we keep any error due to linearization small compared to the errors because of the model mismatch. This allows for confident steps in the parameter update.

### v3

@@ Abstract
We tackle the problem of system identification, where we select inputs, observe the corresponding outputs from the true system, and optimize the parameters of our model to best fit the data. We propose a practical and computationally tractable methodology that is compatible with any system and parametric family of models. Our approach only requires input-output data from the system and first-order information of the model with respect to the parameters. Our approach consists of two modules. First, we formulate the problem of system identification from a Bayesian perspective and use a linear Gaussian model approximation to iteratively optimize the model’s parameters. In each iteration, we propose to use the input-output data to tune the covariance of the linear Gaussian model. This online covariance calibration stabilizes fitting and signals model inaccuracy. Secondly, we define a Gaussian-based uncertainty measure for the model parameters, which we can then minimize with respect to the next selected input. We test our method with linear and nonlinear dynamics.
@@ I Introduction
In this work, we are interested in the problem of fitting a model to a system and choosing informative inputs to improve the model. Our formulation tackles the general problem of model-fitting and dataset collection, which is known as system identification. Obtaining a good model of the system is a key requirement in various engineering tasks. System identification is usually the first step in a larger pipeline that involves the control of the underlying system, e.g., with model predictive control [1]. Fitting model parameters to data can be challenging when the model is a complicated function of the parameters. Ott et al. have shown that selecting control inputs based on the current model can improve the system identification process [2]. Informative input design can reduce the data gathering requirements for system modeling, which can be costly in real-world applications. This can be particularly useful in high-dimensional problems, where finding informative directions to perturb the system can be challenging.
There are four basic elements in the problem of system identification. We need to select a parametric family of models for the underlying system. The family must be expressive enough to capture the true system, but also simple enough to optimize over. Our parameter estimate is improved once more data becomes available. We must also choose the optimization approach that updates the model parameters. A vast literature exists on methods for estimating the model parameters [3]. Furthermore, we need an uncertainty measure to minimize over the next input. This allows us to select inputs that best inform our choice of model parameters. Various objectives based on the Fisher information matrix can be minimized over the input [4]. This optimization problem must also be accompanied by an optimization approach.
@@
  • Can we design a method for system identification that is computationally light, compatible with any parametric model family and able to determine the suitability of the model family for a given system?
Such a method would be useful for practitioners who usually want a fast and model-agnostic approach.
Our approach leverages first-order information to construct a linear Gaussian model for the output given the input, as a function of the model parameters. We use the first-order Taylor expansion with respect to the model parameters and solve successive convex problems to improve the model parameter estimate. This simplifies the model update step discussed in the previous paragraph, which becomes computationally tractable. We also propose a procedure to tune the covariance associated with the linear Gaussian model in an online manner. This allows us to statistically calibrate our model and judge the suitability of the parametric model family. Finally, using the linear Gaussian model we define an uncertainty measure to optimize over the next input.
@@ II Prior Work
The problem of system identification has been well studied. Research on the optimal selection of inputs dates back to the 1960s [5, 6]. Recent advances in computational capabilities and learning-based methods have once again brought attention to the problem [7, 8]. This section summarizes the prior work on parameter estimation and informative input design for system identification.
Estimating the model parameters using input-output data is usually formulated as a regression problem [9]. A regression problem can be equivalently viewed as maximum likelihood estimation (MLE) under an appropriate distribution. For example, least squares regression corresponds to MLE with a Gaussian distribution. When estimating the model parameters, it is common to assume the true system belongs to the parametric model family [3, 4, 10]. Then, under regularity conditions on the distribution of the input or the structure of the system, the parameter estimate is concentrated around the true system parameter [3]. Our work deals with the problem of system identification when the true system does not necessarily belong to the selected parametric model family. In this case, it is critical to obtain a good fit but also a sense of the suitability of the selected model family.
@@
In various applications, the system contains modes that are not properly described by a linear model. In these cases, we must deal with nonlinear models and nonlinear system identification. Early results focused on specific system dynamics for which the input design problem can either be relaxed using linear matrix inequalities [13] or techniques from linear systems can be applied to design informative step inputs [14]. In certain instances, the nonlinearity can be separated from the linear part of the model in what is known as structured nonlinear system identification [4]. There is consensus among researchers that nonlinear system identification is more difficult than linear system identification [15]. This is, to some extent, because nonlinear optimization is more challenging than linear programming. Our approach deals with the identification of nonlinear dynamics using convex programming. Convex programs are solved efficiently using existing packages [16].
We examine methods for nonlinear system identification that are relevant to our approach. Goodwin presented an algorithm for informative input design in the general nonlinear case that is based on linearization of the parametric model [17]. Our work adds to this work in an important way: our method updates the estimate of the model parameters in a way that respects the linearization and is based on convex optimization, which makes it computationally fast. A similar work was done by Babar and Baglietto [18], who propose an extended Kalman filter for the joint state and parameter estimation of the dynamical system. Model predictive control techniques are used to select the next input. We improve upon their work by adding a trust-region constraint to the update of the model parameters after the linearization and by proposing a modified optimization objective for the selection of the next input that does not rely on sampling. In contrast with both works [17, 18], we tune the noise covariance at every iteration. In addition to statistically calibrating our approach, the tuned covariance serves as a lightweight diagnostic about the suitability of the selected model family. We demonstrate this empirically.
@@ III Problem Formulation
where $x\in\mathbb{R}^{d_{x}},y\in\mathbb{R}^{d_{y}}$ and $g:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{y}}$. We refer to $x$ as the input and $y$ as the output. If $g$ represents dynamics, the input is usually the current state and control, which we abstract as $x$, while the next state is abstracted as $y$. We treat system $g$ as a black-box. We can only supply an input $x$ and obtain an output $y$ from the system.
We choose a family of functions $f(\cdot;\theta)$, parameterized by $\theta\in\mathbb{R}^{d_{\theta}}$. At every iteration, our goal is to compute:
@@
[…]
@@
In the present work, we do not deal with the selection process for the parametric family $f(\cdot;\theta)$. Nevertheless, as we will see below, our approach allows the user to test its suitability for the underlying system.
@@ IV Proposed Method
We first outline our Bayesian approach for improving the current estimate of the model parameters. We then use this Bayesian formulation in order to define an optimization problem for the next input. The components presented so far can be found in the literature [17], with the exception of the trust region in the estimate improvement step and the ability to choose the underlying probability density functions. Finally, we discuss the covariance fine-tuning, which is critical in the convergence of the method and constitutes our major contribution.
@@ IV-A Parameter Estimation
[…]
@@
where $\epsilon_{i}$ is the model error for the $i$-th data-point. We treat $\epsilon_{i}$ as a random quantity and choose the underlying model $\epsilon_{i}\sim\mathcal{N}(0,\Sigma)$ for some positive definite covariance $\Sigma$. Tuning this covariance is critical in the performance of the algorithm. Assuming conditionally independent measurements, the posterior for the hidden parameter is given by
@@
where we choose the prior for $\theta$, $p(\theta)$.
The maximum a-posteriori (MAP) estimate for the parameter under these assumptions is given by
@@
$=\operatorname*{arg\;min}_{\theta}\ -\log p(\theta)+\dfrac{1}{2}\sum_{i=1}^{n}\lVert y_{i}-f(x_{i};\theta)\rVert_{\Sigma^{-1}}^{2}.$
In general, this can be a very difficult problem to solve because it depends on the structure of the parametric family $f$ with respect to the parameter $\theta$. We can look for a local optimum using gradient-based techniques. We opt for an approximate treatment of the optimization problem that lends itself to a convex solver [19].
@@
[…]
@@
[…]
@@
$\hat{\theta}_{+}\approx\operatorname*{arg\;min}_{\lVert\theta-\hat{\theta}\rVert\leq\delta}$ $-\log p(\theta)+$ (8)
$\dfrac{1}{2}\sum_{i=1}^{n}\lVert\underbrace{y_{i}-b(x_{i};\hat{\theta})}_{\tilde{y}_{i}}-C(x_{i};\hat{\theta})\theta\rVert_{\Sigma^{-1}}^{2}.$
[…]
We can replace Gaussian densities with any other log concave density and still obtain a convex problem. The used density for $\epsilon_{i}$ can be changed in each estimation step, based on the realized errors $\{y_{i}-f(x_{i};\hat{\theta})\}_{i=1}^{n}$. This allows us to select the family of underlying distributions. More details on this will be given in section IV-C.
@@ IV-B Informative Input Design
After obtaining $\hat{\theta}_{+}$, our goal is to select $x_{n+1}$. The goal of selecting $x_{n+1}$ is to reduce our uncertainty in the posterior estimate of $\theta$. Assuming that we are given a new data-point $(x_{n+1},y_{n+1})$ and under linear Gaussian assumptions (for the prior as well), the posterior inverse covariance is given by the Hessian of the objective of (8) but evaluated using the most up-to-date MAP estimate $\hat{\theta}_{+}$, i.e,
[…]
where we assume $\theta\sim\mathcal{N}(\theta_{\mathrm{prior}},\Sigma_{\mathrm{prior}})$. In the case of a uniform prior for $\theta$, we can set $\Sigma_{\mathrm{prior}}^{-1}=0$, i.e., we can simply ignore the prior term in (9).
It has been shown that the covariance (9) can be accurate in the nonlinear case, although it is exact only in the linear regime [21]. Eq. (9) is an approximate Fisher information matrix, where a prior is added and the expectation over the data is not used. We note that it only depends on the inputs $x_{1},\dots,x_{n+1}$ and not on the corresponding outputs.
@@
[…]
@@
[…]
@@
In order to solve eq. 11 with gradient descent methods, we can replace the indicators with smooth functions. By increasing parameter $\lambda\geq 0$, we can move the search towards feasible inputs $x_{n+1}$. This is an example of a penalty method [20]. As a receding horizon extension, we can easily extend eq. 11 to optimize over the next $k$ inputs, $x_{n+1},\dots,x_{n+k}$. In (11), because the system $g$ is only accessible via a query oracle, we approximate $g(x_{n+1})\in\mathcal{Y}$ by $f(x_{n+1};\hat{\theta}_{+})\in\mathcal{Y}$.
IV-C Fine-tuning
We present the fine-tuning approach for the covariance $\Sigma$ and also propose a way to tune $\delta$.
At the current iteration , we have access to the dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$. After obtaining the new parameter estimate $\hat{\theta}_{+}$, we can compute the estimated noise realizations $\epsilon^{\mathrm{model}}_{i}$ by comparing the true output to the model output, i.e., for $i=1,\dots,n$
[…]
$\Sigma_{\mathrm{model\ error}}=:\dfrac{1}{n}\sum_{i=1}^{n}\epsilon^{\mathrm{model}}_{i}{\epsilon^{\mathrm{model}}_{i}}^{\top}.$ (13)
@@ IV-C Fine-tuning
$y_{i}-b(x_{i};\hat{\theta})-C(x_{i};\hat{\theta})\hat{\theta}_{+}$ (14)
@@
$\epsilon^{\mathrm{model}}_{i}+\underbrace{f(x_{i};\hat{\theta}_{+})-\left(b(x_{i};\hat{\theta})+C(x_{i};\hat{\theta})\hat{\theta}_{+}\right)}_{\epsilon^{\mathrm{lin}}_{i}}.$
This implies that the covariance $\Sigma$ should be set using the sample covariance of $\{\epsilon^{\mathrm{model}}_{i}+\epsilon^{\mathrm{lin}}_{i}\}_{i=1}^{n}$:
$\Sigma\leftarrow\dfrac{1}{n}\sum_{i=1}^{n}\left(\epsilon^{\mathrm{model}}_{i}+\epsilon^{\mathrm{lin}}_{i}\right)\left(\epsilon^{\mathrm{model}}_{i}+\epsilon^{\mathrm{lin}}_{i}\right)^{\top},$ (16)
where $\epsilon^{\mathrm{lin}}_{i}$ is the error between the value from the model and the value from the Taylor expansion of the model for the $i$-th output.
For setting the trust region $\delta$, we can use a simple heuristic. If the average linearization error size $n^{-1}\sum_{i=1}^{n}\lVert\epsilon^{\mathrm{lin}}_{i}\rVert$ becomes comparable to the average absolute model error size $n^{-1}\sum_{i=1}^{n}\lVert\epsilon^{\mathrm{model}}_{i}\rVert$, we reduce $\delta$. This procedure guarantees that we keep any error due to linearization small compared to the errors because of the model mismatch. This allows for confident steps in the parameter update.